# MSCCL++: Rethinking GPU Communication Abstractions for Cutting-edge AI Applications

### v1

@@ 2.3 Limitations of Existing User Interfaces
NCCL API is designed to provide end-to-end implementations of collective communication operations. This would be what most applications expect from the library, but this is fundamentally limited to support diverse performance-sensitive applications. Like other GPU workloads do, collective communication over GPUs should be differently optimized for different scenarios, depending on the input size, GPU architecture, networking topology, and the application characteristics [32, 12]. It is very difficult for the library developers to provide the best performance for all cases in a timely manner.
@@
MSCCL [9] is a previous work that tackles this issue by allowing developers to implement their own custom communication algorithms using a DSL. However, MSCCL is still based on the NCCL primitives and does not provide enough flexibility to fully utilize the hardware resources, as described in Section 2.2.2. The MSCCL++ Primitive API we propose in this paper is designed at a lower level that barely abstracts the GPU hardware, which enables optimization in the hardware’s perspective. This would provide a more flexible and efficient way to optimize GPU communication. Inspired by MSCCL, we also provide a DSL for software developers to easily develop communication algorithms.
3 MSCCL++ Overview
Figure 3: Overview of MSCCL++.
3.1 Hierarchical User Interfaces
@@
MSCCL++ provides three user APIs at hierarchical levels of communication abstractions, as shown in Figure 3: Primitive, DSL, and Collective APIs. The three levels provide different trade-offs in terms of programming effort or expertise and execution latency or performance. The closer the interface is to the hardware, the higher the programmer’s control on execution, so higher the performance and programming effort.
@@
MSCCL++ Primitive API (or MSCCL++ API). This replaces NCCL’s primitive API with a set of core communication functions that are called from the GPU kernels. This is paired with the bootstrapping API called from the host (CPU) side, which is used for connection setup between GPUs. The primitive interface is the base layer that is used to implement the other two higher-level interfaces. Programmers can also use the primitive interface directly to implement application-specific or hardware-specific optimizations in their own GPU kernels, including those kernels that also perform computation. As the primitive interface is our key contribution, we focus on its details throughout Section 3 and 4.
@@
MSCCL++ DSL API. We reimplement the MSCCLang [10] DSL over MSCCL++, which allows to write custom collective communication algorithms in a high-level language. This interface is aimed at users who want to generate communication kernels that are optimized for their own workloads (application and hardware). We elaborate in Section 4.3.
@@
MSCCL++ Collective API. We reimplement the NCCL API, including its bootstrapping API, as-is over MSCCL++. It is aimed at users with the least expertise; they can simply replace NCCL/RCCL with the MSCCL++ Collective library without changing their application code. This library has the same limitations as that of NCCL (Section 2.3); i.e., it may not provide the best algorithm for certain workloads. For better performance, users can install their own optimized algorithms written using the MSCCL++ DSL API into this library.
3.2 Primitive Communication Abstractions
3.2.1 Communication Channels
@@
Since the mode(s) of data transfer supported differs between interconnects, MSCCL++ defines an abstract channel for each data transfer mode supported by the hardware and exposes communication primitives specific to that channel that can be invoked directly from inside a GPU kernel. MSCCL++ defines three communication channels [^7] For readers who are familiar with NCCL: MSCCL++ channel does not correspond to that in NCCL.: PortChannel, MemoryChannel, and SwitchChannel, which correspond to port-mapped I/O, memory-mapped I/O, and switch-mapped I/O respectively. A PortChannel uses interconnect ports to communicate between GPUs; i.e., a GPU can initiate the data transfer through a port to another GPU (note that ports are controlled by dedicated hardware for I/O, such as DMA engines on GPUs or RDMA NICs). A MemoryChannel uses peer-to-peer memory access to communicate between GPUs; i.e., a GPU can directly access another GPU’s memory. A SwitchChannel uses interconnection-switch-enabled multimem memory access to communicate between GPUs; i.e., a GPU can access the memory of multiple GPUs simultaneously through a switch.
@@
The same interconnect can support multiple modes of data transfer, so multiple channels can be supported on the same interconnect; for example, all three channels are supported by NVLink, only PortChannel and MemoryChannel are supported by xGMI and PCIe, while only PortChannel is supported by InfiniBand. The implementation of each channel is specific to the interconnect. The SwitchChannel uses NVSwitch multimem instructions that support computation-on-switch. The MemoryChannel uses thread-copy data transfer mode over NVLink, xGMI, and PCIe. On the other hand, the PortChannel uses the DMA-copy data transfer mode over NVLink, xGMI, PCIe, and Infiniband. Current hardware interconnects require a CPU thread to initiate the data transfer. For example, transferring data over InfiniBand requires the CPU to call ibv_post_send that initiates the RDMA transfer from the sending GPU memory to the receiving GPU memory. However, such details are handled transparently by the implementation.
@@
Figure 4: MSCCL++ data transfer abstractions. put asynchronously transfers data from one GPU to another. signal and wait synchronize data transfer between GPUs. flush ensures the completion of preceded data transfer.
3.2.2 Communication Primitives
Borrowing the term from MPI [34], writing data to the peer’s side through a channel is called put. To provide our communication abstraction as close as possible to hardware capabilities, we design put to be zero-copy (i.e., no intermediate buffers), one-sided (i.e., initiated by a peer without participation of the other peer), and asynchronous. MSCCL++ provides other interfaces for synchronization purposes. Specifically, signal and wait provide a mechanism for synchronization (including memory consistency) across GPUs. flush is for local synchronization that ensures all previous put operations are already on the fly.
@@ 3.2.2 Communication Primitives
12 //Reduce ever pair of GPU
@@
Figure 5: All-pairs ReduceScatter kernel in MSCCL++. Channels are initialized with source as input and destination as scratch buffer.
Example. Figure 5 shows MSCCL++ implementation of the all-pairs ReduceScatter algorithm, which is not used in NCCL. In this algorithm, all GPUs send their data at $\frac{i}{N}$ offset to GPU $i$, the $i^{th}$ GPU does elementwise reduction of the received data, and stores the reduced data. To store the received data, the implementation allocates a scratch buffer per GPU. The kernel’s main loop performs reduction in batches of the scratch size (line 518). On each GPU, the loop first put $\frac{1}{N}$ of data to all other GPUs’ scratch buffer and then signals the completion of transfer (line 910). Then a GPU waits for the data to arrive from other GPUs and reduce it with its part of input data (line 1315). Finally, a barrier among all GPUs is required to ensure scratch is not overwritten for next main loop offset (line 1818). Section 5 shows that this algorithm works better than the ring algorithm for smaller data sizes while the ring algorithm works better for large data sizes.
3.3 Advantages
The MSCCL++ Primitive API itself does not introduce any novel optimization techniques – instead, it offers new straightforward abstractions that enable quick optimization of GPU communication. The key advantage comes from exposing the primitive functionalities as a user interface, which introduces three benefits. First, it allows users to easily customize GPU communication algorithms for their own infrastructure and workloads (see following paragraphs). This is valuable in practice because the library developers may not be aware of the specific workloads and hardware configurations of their users, hence the difficulties of timely support of efficient implementations. Second, the library developers can quickly deliver new hardware support by leaving the algorithm design to the users who can do it better (see Section 4.5.1). Third, since the primitive interface is close to the hardware, it is more flexible to adopt new hardware features.
We carefully optimized GPU communication in the MSCCL++ DSL Executor and Collective kernels by leveraging the features in the MSCCL++ Primitive API. Users of the Primitive API can also implement these optimizations directly. Section 5 shows how these optimizations help MSCCL++ obtains better performance than NCCL.
Asynchronous Communication. MSCCL++ contains separate primitives for data transfer and synchronization, thus, enabling asynchronous communication and batching the synchronization of multiple communication into a single synchronization. This abstraction has three benefits: (i) allows more algorithms to be specified than the synchronous communication in NCCL, (ii) enables faster collective communication performance by allowing an algorithm space that can batch several synchronizations into one, (iii) frees up resources of peer GPUs to do other meaningful work inside the application, and (iv) allows better inter-kernel optimizations.
@@ 3.3 Advantages
Computation-Communication Optimizations. Since MSCCL++ Primitive API is an in-kernel interface for communication, we can co-optimize both computation and the communication kernels, leading to better application performance. One such optimization is fusing communication with a GeMM kernel [15, 37, 17, 8, 39], so that the number of memory reads/writes decreases, and a GPU’s warp context switching enables overlapping of communication with the GeMM kernel. MSCCL++ Primitive API is an easy-to-use and efficient interface to implement this idea, and we remain the implementation as a future work.
4 Implementation
4.1 Initialization
In general, an MSCCL++ program runs as a multi-process on all nodes and GPUs within a node, such that each process is responsible for one GPU. Since MSCCL++ enables application-specific optimization, it provides CPU side interfaces for custom initialization of the collective communications used in an application. MSCCL++ provides a default initialization process that can be specialized by users for their application. The initialization process involves setting up connections between GPUs, finding the topology of GPU connections, and constructing channels and sharing scratch buffers between peer GPUs.
The first step in initialization is to create a communication channel between all distributed CPU processes to exchange metadata between processes. We call this communication channel a Bootstrap. A bootstrap consists of four virtual methods: (i) send to send data from one CPU process to another, (ii) recv to recv data from a CPU process, (iii) allGather to perform the AllGather collective communication, and (iv) barrier to synchronize all distributed processes. MSCCL++ implements the default bootstrap methods using POSIX sockets. If a user prefers another distributed communication protocol such as MPI or torch.distributed, it is straightforward to override bootstrap methods because all these methods directly correspond to existing distributed protocols.
The second step for application is to construct a communicator object using the bootstrap object. The application uses the communicator object to register buffers that are used to perform data transfers. These buffers can be either application buffers themselves or scratch buffers dedicated for communications. The communicator then creates relevant channel objects (PortChannel, MemoryChannel, or SwitchChannel) between GPUs based on physical links. For each GPU, these channel objects are initialized with buffers. Finally, these channels are shared among processes of each GPU.
4.2 Primitives for Channels and Protocols
A connection between GPUs in MSCCL++ is called a channel. PortChannel and MemoryChannel are peer-to-peer connections (i.e., between 2 GPUs), whereas SwitchChannel is a connection among a group/collection of (2 or more) GPUs. Figure 6 shows three types of channels and its interface in MSCCL++. A channel is initialized with the source and destination buffers, and an integer allocated on the GPU memory that servers as semaphore. In this section, we describe all three kinds of MSCCL++ channels.
@@ 22 T read(ulong off, uint tid, uint flag=0)
Figure 6: PortChannel, MemoryChannel, and SwitchChannel primitives in MSCCL++. Implementation of MemoryChannel primitives differ for the type of protocol: LL or HB.
4.2.1 PortChannel
A PortChannel implements primitives when data transfer is done over ports connected to GPU memory, such as using InfiniBand API’s ibv_post_send for IB or cudaMemcpy for DMA copy. Since data transfer over a port require the CPU to initiate the transfer, this channel creates a CPU thread for each GPU. The CPU thread reads data transfer or synchronization requests from a first-in-first-out request queue. The storage, head, and tail of a request queue are allocated using cudaMallocManaged, so that both CPU and GPU can access this data. By default we allocate a queue of 1024 requests but this size can be changed by the user. We now discuss the workflow of PortChannels using Figure 7.
Figure 7: PortChannel workflow for IB starting from ⓪ when GPU 0 calls put primitive to ⑧ when GPU 1 receives the data.
0
When the GPU calls a primitive, such as, put, then the first participating thread of the GPU pushes the request to the queue by writing at the head. Before writing to the queue, the GPU checks if the queue is filled, i.e., if the head value is more than the tail value. If the queue is filled then the GPU waits for the CPU to process atleast one request. 1 Then the GPU increments the head to the next element. 2 The CPU thread continously reads the element at the tail to see if there is a request from the GPU and when there is a request, then the CPU reads the request, zeros out the current element, and increments the tail. 3 Now the CPU thread will process the request. Below we explain the requests generated by primitives and how the CPU handle these requests.
@@ 4.2.1 PortChannel
Flush. Primitive flush pushes a flush request in the queue and the first thread of the GPU wait until the queue head is equal to or more than the queue tail. The CPU process the flush request by waiting until all the previous data transfer and synchronization requests have been completed. For example, for IB we use ibv_poll_cq to get the status of all requests. After the flush request is complete, the GPU is free to re-write to the source buffer.
4.2.2 MemoryChannel
@@
(i) In HB protocol, peer GPUs transfer a large chunk of data and synchronize this chunk once using wait and signal primitives. Hence, the synchronization time is amortized over the transfer time leading to high bandwidth. However, since the receiving GPU needs to wait for the whole chunk before the GPU can process it, the HB protocol also has high latency.
@@
(ii) In LL protocol, the receiving GPU synchronizes on fewer chunks of data being transferred and process the chunk as soon as it is received, thus, achieving lower latency. However, since the number of synchronization is proportional to the number of elements, this protocol provides lower-bandwidth. The LL protocol works as follows. The put primitive requires an extra integer flag value. For every $N-1$ elements written to the receiving GPU, put also write the flag. The receiving GPU uses the read primitive waits until the flag value at $N$ index of receiving buffer is set and then read and return the $N-1$ elements. However, we cannot arbitrarily use any $N$ because GPUs follow a weak memory consistency model where writes to different memory locations by multiple threads can be performed in any order. Therefore, we restrict $N$ to number of elements written by a single instruction, i.e., 4, 8, and 16 bytes memory accesses instructions. The flag value is decided based on the algorithm, such that, all flag values are distinct.
@@
Implementation. A MemoryChannel is initialized with a protocol. A protocol defines the implementation of data transfer and synchronization primitives. We now describe the implementation of primitives based on protocol as follows:
@@
Data Transfer. Primitive put reads data elements from source buffer and writes to destination buffer starting from different offsets. Similarly, the read primitive reads and return data from the destination buffer, and the write primitive writes data including the flag for the LL protocol to the destination buffer. To maximize bandwidth for the HB protocol, both primitives use 16-bytes loads and stores. For the LL protocol, both primitives also takes a flag and uses 8-byte load and stores by default or user supplied vector length. Since put both primitives are called by multiple threads (all or first few threads of the kernel), we achieve maximum bandwidth.
@@
Synchronization. Synchronization primitives wait on a semaphore which is an integer allocated on the receiving GPU. The signal primitive atomically increments the semaphore of the receiving GPU and calls threadfence_system to ensure that writes by put and semaphore increments are made available in this order. The wait primitive performs a busy-wait while-loop that checks until the value of semaphore has reached the expected value. This wait is performed by the first thread of the kernel and all other threads wait on a kernel barrier. The channel tracks the expected value using its expectedValue member. Since after put returns memory writes are already in operation, src buffer can be reused. Therefore, the flush primitive is empty.
4.2.3 SwitchChannel

### v2

@@ 2.3. Existing Communication Libraries
NCCL API is designed to provide end-to-end implementations of collective communication operations. This would be what most applications expect from the library, but this is fundamentally limited to support diverse performance-sensitive applications. Like other GPU workloads do, collective communication over GPUs should be differently optimized for different scenarios, depending on the input size, GPU architecture, networking topology, and the application characteristics (Shah et al., 2023; Hwang et al., 2023a). It is very difficult for the library developers to provide the best performance for all cases in a timely manner.
@@
MSCCL (Cowan et al., 2022) tackles this issue by allowing developers to implement their own custom communication algorithms using a DSL (Cowan et al., 2023). However, MSCCL is still based on the NCCL primitives and does not provide enough flexibility to fully utilize the hardware resources, as described in Section 2.2.2.
3. MSCCL++ Overview
Figure 3. Overview of the MSCCL++ stack.
3.1. Hierarchical User Interfaces
@@
MSCCL++ provides three user APIs at hierarchical levels of communication abstractions: Primitive, DSL, and Collective APIs (see Figure 3). The three levels provide different trade-offs in terms of programming effort or expertise and execution latency or performance. The closer the interface is to the hardware, the higher the programmer’s control on execution, so higher the performance and programming effort.
@@
MSCCL++ Primitive API. This replaces NCCL’s primitives with a set of core communication functions that are called from the GPU kernels. This is paired with the bootstrapping API called from the host (CPU) side, which is used for connection setup between GPUs. The primitive interface is the base layer that is used to implement the other two higher-level interfaces. Programmers can also use the primitive interface directly to implement application-specific or hardware-specific optimizations in their own GPU kernels, including those kernels that also perform computation. As the primitive interface is our key contribution, we focus on its details in Section 3 and 4.
@@
MSCCL++ DSL API. We reimplement the MSCCLang (Cowan et al., 2023) DSL over MSCCL++, which allows us to write custom collective communication algorithms in a high-level language. This interface is aimed at users who want to specify communication algorithms or schedules and generate kernels that are optimized for their own workloads (application and hardware). The generated kernel uses the Primitive API by running the DSL Executor, so it may have small runtime overheads.
@@
MSCCL++ Collective API. We reimplement the NCCL API, including its bootstrapping API, as-is over MSCCL++. It is aimed at users with the least expertise; they can simply replace NCCL/RCCL with the MSCCL++ Collective library without changing their application code. This library has the same limitations as that of NCCL (Section 2.3); i.e., it may not provide the best algorithm for certain workloads. For AI inference workloads, we implement the best algorithms in our collective kernels using the SpeCL Primitive API.
3.2. Primitive Communication Abstractions
3.2.1. Communication Channels
@@
The supported mode(s) of data transfer differs between interconnects. Therefore, MSCCL++ defines an abstract channel for each data transfer mode supported by the hardware, and exposes communication primitives specific to that channel that can be invoked directly inside a GPU kernel. MSCCL++ defines three communication channels (for readers who are familiar with NCCL, MSCCL++ channel does not correspond to that in NCCL): PortChannel, MemoryChannel, and SwitchChannel, which correspond to port-mapped I/O, memory-mapped I/O, and switch-mapped I/O respectively. A PortChannel uses interconnect ports to communicate between GPUs; i.e., a GPU can initiate the data transfer through a port to another GPU. Ports are controlled by dedicated hardware for I/O, such as DMA engines on GPUs or RDMA NICs. A MemoryChannel uses peer-to-peer memory access to communicate between GPUs; i.e., a GPU can directly access another GPU’s memory. A SwitchChannel uses interconnection-switch-enabled multimem memory access to communicate between GPUs; i.e., a GPU can access the memory of multiple GPUs simultaneously through a switch.
@@
The same interconnect can support multiple modes of data transfer, so multiple channels can be supported on the same interconnect; e.g., all three channels are supported by NVLink, only PortChannel and MemoryChannel are supported by xGMI and PCIe, while only PortChannel is supported by InfiniBand. The implementation of each channel is specific to the interconnect. The SwitchChannel uses NVSwitch multimem instructions that support computation-on-switch. The MemoryChannel uses thread-copy data transfer mode over NVLink, xGMI, and PCIe. On the other hand, the PortChannel uses the DMA-copy data transfer mode over NVLink, xGMI, PCIe, and Infiniband. Current interconnects require a CPU thread to initiate the data transfer; e.g., transferring data over InfiniBand requires the CPU to call ibv_post_send which initiates the RDMA transfer from the sending GPU memory to the receiving GPU memory. Such details are handled transparently by the implementation.
@@
3.2.2. Communication Primitives
Borrowing the term from MPI (Snir et al., 1998), writing data to the peer’s side through a channel is called put. To provide our communication abstraction as close as possible to hardware capabilities, we design put to be zero-copy (i.e., no intermediate buffers), one-sided (i.e., initiated by a peer without participation of the other peer), and asynchronous. MSCCL++ provides other interfaces for synchronization purposes. Specifically, signal and wait provide a mechanism for synchronization (including memory consistency) across GPUs. flush is for local synchronization that ensures all previous put operations are already on the fly.
@@ 3.2.2. Communication Primitives
Figure 4. MSCCL++ data transfer abstractions. put asynchronously transfers data from one GPU to another. signal and wait synchronize data transfer between GPUs. flush ensures the completion of preceded data transfer.
Example. Figure 5 shows MSCCL++ implementation of the all-pairs ReduceScatter algorithm, which is not used in NCCL. In this algorithm, all GPUs send their data at $\frac{i}{N}$ offset to GPU $i$, the $i^{th}$ GPU does elementwise reduction of the received data, and stores the reduced data. To store the received data, the implementation allocates a scratch buffer per GPU. The kernel’s main loop performs reduction in batches of the scratch size (line 518). On each GPU, the loop first put $\frac{1}{N}$ of data to all other GPUs’ scratch buffer and then signals the completion of transfer (line 910). Then a GPU waits for the data to arrive from other GPUs and reduce it with its part of input data (line 1315). Finally, a barrier among all GPUs is required to ensure scratch is not overwritten for next main loop offset (line 18). Section 5 shows that this algorithm works better than the ring algorithm for smaller data sizes while the ring algorithm works better for large data sizes.
3.2.3. Advantages of Communication Primitives
@@
12 //Reduce every pair of GPU
@@
Figure 5. All-pairs ReduceScatter kernel in MSCCL++. Channels are initialized with source as input and destination as scratch buffer.
The MSCCL++ Primitive API itself does not introduce any novel optimization techniques – instead, it offers new straightforward abstractions that enable quick optimization of GPU communication. The key advantages come from exposing the primitive functionalities as a user interface. First, it allows application developers to easily and quickly optimize GPU communication algorithms for their own infrastructure and workloads. This is valuable in practice because the library developers may not be aware of the specific workloads and hardware configurations of their users. Second, the library developers can quickly deliver new hardware support by implementing the same primitive API for the new hardware. Third, since the primitive API is close to the hardware, it is more flexible to extend or adopt it for new hardware features.
MSCCL++ carefully optimizes GPU communication in the DSL Executor and collective kernels leveraging some of the below features in the Primitive API. Users of the Primitive API can also implement these optimizations directly.
Asynchronous Communication. MSCCL++ contains separate primitives for data transfer and synchronization, allowing asynchronous communication and batching synchronization of multiple data transfers. This abstraction has multiple benefits: (i) allows more algorithms to be specified than the synchronous communication in NCCL, (ii) enables faster collective communication performance by enabling algorithms to combine several synchronizations into one, (iii) frees up resources of peer GPUs for other meaningful work of the application, and (iv) enables better inter-kernel optimizations.
@@ 22 T read(ulong off, uint tid, uint flag=0)
Figure 6. PortChannel, MemoryChannel, and SwitchChannel primitives in MSCCL++. Implementation of MemoryChannel primitives differ for the type of protocol: LL or HB.
4. SpeCL Design and Implementation
4.1. Initialization of SpeCL
In general, a MSCCL++ program runs as a multi-process on all nodes and GPUs within each node, so that each process runs one GPU. MSCCL++ provides host (CPU) side interfaces for initializing multiple processes, and the interfaces can be specialized by users for their application. The initialization involves setting up connections between GPUs, finding the topology of GPU connections, and constructing channels and buffers.
The first step in initialization is to exchange metadata between host processes, which we call a Bootstrap. It consists of four virtual methods: (i) send to send metadata to another process, (ii) recv to receive metadata from a process, (iii) allGather to collect metadata of all processes, and (iv) barrier to synchronize all processes. MSCCL++ implements a default bootstrap methods using POSIX sockets, while it is straightforward to override with other popular implementations such as MPI or torch.distributed.
The second step is to construct a Communicator that registers communication buffers (either user buffers or dedicated ones) and creates corresponding MSCCL++ channels between GPUs based on the underlying physical links.
4.2. MSCCL++ Primitive API and Channels
A MSCCL++ channel is a connection between two or more GPUs. PortChannel and MemoryChannel are peer-to-peer (i.e., between two GPUs), whereas SwitchChannel is for a group of (i.e., two or more) GPUs. Figure 6 shows three types of channels and its interface in MSCCL++. A channel is initialized with the source and destination buffers, and an integer allocated on the GPU memory that servers as a semaphore.
4.2.1. PortChannel
A PortChannel implements primitives when data transfer is done over ports connected to GPU memory, such as cudaMemcpy for intra-node DMA or ibv_post_send for RDMA. Since data transfer over a port currently requires the CPU to initiate the transfer, each channel creates its own CPU thread. The CPU thread reads data transfer or synchronization requests from a first-in-first-out request queue. The storage, head, and tail of a request queue are allocated using cudaMallocManaged as both CPU and GPU can access it.
We now discuss the workflow of PortChannels using Figure 7. 0 When the GPU calls a primitive, such as put, the first participating thread of the GPU pushes the request to the queue by writing at the head. Before writing to the queue, the GPU checks if the queue is filled, i.e., if the head value is more than the tail value. If the queue is filled then the GPU waits for the CPU to process atleast one request. 1 Then the GPU increments the head to the next element. 2 The CPU thread continuously reads the element at the tail to see if there is a request from the GPU, and when there is a request, the CPU reads the request, zeros out the current element, and increments the tail. 3 Now the CPU thread will process the request. Below we explain the requests generated by primitives and how the CPU handle these requests.
Figure 7. PortChannel workflow for IB from ⓪ when GPU 0 calls put primitive to ⑧ when GPU 1 receives the data.
@@ 4.2.1. PortChannel
Flush. Primitive flush pushes a flush request in the queue and the first thread of the GPU waits until the queue head is equal to or more than the queue tail. The CPU processes the flush request by waiting until all the previous data transfer and synchronization requests have been completed. For example, for IB we use ibv_poll_cq to get the status of all requests. After the flush request is complete, the GPU is free to re-write to the source buffer.
4.2.2. MemoryChannel
@@
(i) In the HB protocol, peer GPUs transfer a large chunk of data and synchronize this chunk once using wait and signal primitives. Hence, the synchronization time is amortized over the transfer time leading to high bandwidth. However, since the receiving GPU needs to wait for the whole chunk before processing it, the HB protocol also has high latency.
@@
(ii) In LL protocol, the receiving GPU synchronizes on fewer chunks of data being transferred and process the chunk as soon as it is received, thus, achieving lower latency. However, since the number of synchronization is proportional to the number of elements, this protocol provides lower-bandwidth. The LL protocol works as follows. The put primitive requires an extra integer flag value. For every $N-1$ elements written to the receiving GPU, put also writes the flag. The receiving GPU uses the read primitive, waits until the flag value at $N$ index of receiving buffer is set, and then reads and returns the $N-1$ elements. However, we cannot arbitrarily use any $N$ because GPUs follow a weak memory consistency model where writes to different memory locations by multiple threads can be performed in any order. Therefore, we restrict $N$ to number of elements written by a single instruction, i.e., 4, 8, and 16 bytes memory accesses instructions. The flag value is decided based on the algorithm, such that, all flag values are distinct.
@@
Implementation. A MemoryChannel is initialized with a protocol. A protocol defines the implementation of data transfer and synchronization primitives. We now describe the implementation of primitives based on the protocol as follows:
@@
Data Transfer. Primitive put reads data elements from source buffer and writes to destination buffer starting from different offsets. Similarly, the read primitive reads and returns data from the destination buffer, and the write primitive writes data including the flag for the LL protocol to the destination buffer. To maximize bandwidth for the HB protocol, both primitives use 16-bytes loads and stores. For the LL protocol, both primitives also takes a flag and uses 8-byte load and stores by default or user supplied vector length. Since both put primitives are called by multiple threads (all or first few threads of the kernel), we achieve maximum bandwidth.
@@
Synchronization. Synchronization primitives wait on a semaphore that is an integer allocated on the receiving GPU. The signal primitive atomically increments the receiving GPU’s semaphore and calls threadfence_system to ensure that writes by put and semaphore increments are made available in this order. The wait primitive performs a busy-wait while-loop that checks until the value of semaphore has reached the expected value. This wait is performed by the first thread of the kernel and all other threads wait on a kernel barrier. The channel tracks the expected value using its expectedValue member. The flush primitive is a no-op because after put returns, the source buffer can be reused, even though the write is still in progress.
4.2.3. SwitchChannel